Alycia opens her piggy bank to find only snickels and dimes inside. She counts the coins and finds that she has 5757 coins totaling ​$4.85 How many nickels were in the piggy​ bank, and how many dimes were in​ there?
Found 2 solutions by ikleyn, Alan3354:
Answer by ikleyn(52797)

.
Wrong data.


Answer by Alan3354(69443)

Alycia opens her piggy bank to find only snickels and dimes inside. She counts the coins and finds that she has 5757 coins totaling $4.85 How many nickels were in the piggy bank, and how many dimes were in there?
----------
If it's 57 coins:
n + d = 57
5n + 10d = 485 --> n + 2d = 97
------
n + d = 57
n + 2d= 97
------------------ Subtract